OpenStudy (anonymous):

Some birds and squirrels were on a branch of a tree. There were 15 animals on the branch with a total of 50 animal feet. How many birds and how many squirrels were on the branch? Part 1: Set up the problem. (2 points) Part 2: Explain why you set it up that way. (1 point) Part 3: Provide the solution. (1 point)

OpenStudy (anonymous):

part 1: let b = number of birds... let s = number of squirrels.... Then, 2b = number of bird feet.. 4s = number of squirrel feet.. So, part 2: your 2 equations will be: b + s = 15 (animals) 2b + 4s = 50 (animal feet)

OpenStudy (anonymous):

Then, to solve, use substitution: b = 15 - s 2b + 4s = 50.. so, 2(15 - s) + 4s = 50 30 - 2s + 4s = 50 s = 10 (10 squirrels)... and, then 5 birds!!